# Universal Fine-Grained Symmetry Inference and Enforcement for Rigorous Crystal Structure Prediction

### v1

@@ Abstract
Crystal structure prediction (CSP), which aims to predict the three-dimensional atomic arrangement of a crystal from its composition, is central to materials discovery and mechanistic understanding. Existing deep learning models often treat crystallographic symmetry only as a soft heuristic or rely on space group and Wyckoff templates retrieved from known structures, which limits both physical fidelity and the ability to discover genuinely new material structures. In contrast to retrieval-based methods, our approach leverages large language models to encode chemical semantics and directly generate fine-grained Wyckoff patterns from composition, effectively circumventing the limitations inherent to database lookups. Crucially, we incorporate domain knowledge into the generative process through an efficient constrained-optimization search that rigorously enforces algebraic consistency between site multiplicities and atomic stoichiometry. By integrating this symmetry-consistent template into a diffusion backbone, our approach constrains the stochastic generative trajectory to a physically valid geometric manifold. This framework achieves state-of-the-art performance across stability, uniqueness, and novelty (SUN) benchmarks, alongside superior matching performance, thereby establishing a new paradigm for the rigorous exploration of targeted crystallographic space. This framework enables efficient expansion into previously uncharted materials space, eliminating reliance on existing databases or a priori structural knowledge.
@@
Crystalline materials are fundamental to modern technologies spanning energy, electronics, medicine, and aerospace, and advancing these sectors relies on identifying novel crystal structures with tailored properties [7, 3, 10, 6, 15]. The vastness of the compositional and structural parameter space presents a fundamental bottleneck for materials discovery. Conventional pipelines, guided by chemical intuition and constrained by costly computations and experiments [21, 16], are intrinsically inefficient at exploring this space at scale. In recent years, artificial intelligence (AI), and deep learning-based generative models in particular, has emerged as a promising route to proposing candidate materials or predicting their properties at scale, demonstrating superior generalization capabilities and significant advantages in computational efficiency [11, 2, 13, 28, 31, 8, 30, 29]. Within this landscape, the crystal structure prediction (CSP) task [11, 13, 8], which seeks the three-dimensional arrangement of atoms in the unit cell from a given composition, plays a pivotal role in AI-driven materials design workflows [28]. Diffusion models have become the dominant architecture for 3D structure generation [26, 24], and, for the CSP task, diffusion-based approaches also deliver state-of-the-art performance [13, 30].
However, the CSP task imposes particularly stringent requirements on the design of AI methods. Physically plausible crystal structures typically obey the symmetry constraints imposed by specific space groups. Learning and explicitly enforcing these symmetries can substantially reduce the search space of CSP, thereby improving both the efficiency and success rate of structure generation. Crystal symmetry strictly constrains the allowed Wyckoff positions, determining atomic multiplicities, site symmetries, and the relative arrangement of atoms within the unit cell. Only when the given chemical composition is mathematically consistent with the corresponding Wyckoff positions can physically reasonable crystal structures be obtained.
@@
However, existing AI methods only partially address this requirement. Some recent approaches on generating atomic structures condition the generator on a space group label [28, 30] but do not explicitly represent or enforce Wyckoff positions, so symmetry is only captured in a coarse global sense. Other works attempt to encode Wyckoff templates into a latent representation that modulates the generative process [31, 4], yet their modulation mechanisms do not guarantee that the final atomic coordinates satisfy the exact Wyckoff constraints. DiffCSP++ [13] goes further by enforcing strict space group symmetry defined by Wyckoff templates, but it assumes that a suitable space group and Wyckoff template are already known in the given lookup database. In practice, these templates are retrieved from existing structures via metric learning methods, such as CSPML [14]. However, restricting generation to retrieved templates confines the model to known symmetry patterns, thereby limiting its ability to discover genuinely new space-group and Wyckoff configurations for a given composition.
@@
To address these fundamental limitations, we propose a symmetry-driven generative framework, which enables the ab initio generation of fine-grained Wyckoff site assignments directly from composition and atom counts, independent of any structural priors, as depicted in Fig. 1. Crucially, the predicted symmetries are imposed as hard geometric constraints to guide and modulate the diffusion process for three-dimensional crystal geometries, thereby rectifying the denoising trajectory. By ensuring strict compliance with Wyckoff symmetry, our approach confines the generative process to a physically valid manifold, guaranteeing both structural plausibility and computational precision. As a result, this leads to a significant enhancement in the stability, uniqueness, and novelty (SUN) [30] of the discovered materials, as well as the matching rate metrics [12].
@@
Below, we outline the core methodology, with full implementation details and training configurations provided in Supplementary Information A and B, respectively.
@@
The second LLM, denoted as $LLM_{w}$, takes the same expanded atomic sequence together with the predicted space group $S_{g}$ and outputs a probability distribution over Wyckoff letters compatible with that space group for each atom. Rather than concatenating $S_{g}$ naively to the input, we encode the predicted space group into a continuous embedding and inject it into each transformer block via the feature-wise linear modulation (FiLM) [18] module. This FiLM-based conditioning ensures that Wyckoff predictions are explicitly modulated by the inferred space group, rather than being learned in an implicitly entangled fashion. Concretely, FiLM($S_{g}$) produces per-channel scaling and shifting coefficients that modulate the intermediate representations of $LLM_{w}$, allowing the model to adapt its representation to the predicted symmetry class:
$p_{w}(l_{n}|C_{\text{atoms}})=LLM_{w}\big(C_{\text{atoms}},\mathrm{FiLM}(S_{g})\big),$ (2)
@@
where $p_{w}(l_{n}|C_{\text{atoms}})$ denotes the probability that the $n$-th atom is assigned to letter $l$, $L(S_{g})$ is the set of admissible Wyckoff letters for space group $S_{g}$. Specific details regarding the network architectures and hyperparameter configurations of these two large language models are provided in Supplementary Information A.1-A.2 and B.1-B.2.
Second, we formulate the Wyckoff letter assignment for each atom as a constrained optimization problem based on the probability distribution $p_{w}(l_{n}|C_{\text{atoms}})$ predicted by the LLM. This formulation strictly adheres to the combinatorial constraints imposed by the Wyckoff site multiplicity rules under the inferred space group symmetry. Specifically, the constrained maximization objective is defined as follows:
@@
where $\text{mult}(l_{n})$ refers to the multiplicity of Wyckoff letter $l_{n}$, and $\text{Count}_{e}(l_{n})$ is the number of atoms of element $e$ assigned to $l_{n}$, $N_{e}$ is the number of atoms of element $e$ in a cell, the set $\{l_{n}^{*}\}_{n=1}^{N}$ represents the optimal Wyckoff letter assignment. The key point here is that $\text{Count}_{e}(l_{n})$ must be an integer multiple of $\text{mult}(l_{n})$, a hard constraint that cannot be satisfied by simply using the maximum probability outputs of each token from the LLM. However, finding the solution for Eq. (3) via brute-force enumeration is computationally intractable, as the search space grows exponentially with the number of atoms, scaling as $O(|L(S_{g})|^{N})$. To overcome this bottleneck, we employ the constrained beam search algorithm to efficiently solve the optimization problem. By maintaining only the most promising candidates within a fixed beam width, this strategy drastically compresses the time complexity from exponential to a linear $O(N)$, rendering the prediction of physically valid Wyckoff assignments computationally tractable. Comprehensive details regarding this algorithm are provided in Supplementary Information A.3.
Third, we guide and constrain the diffusion-based generation of 3D atomic structures using the predicted space group and Wyckoff site template. To instantiate this approach, we adopt the graph neural network (GNN) applied by DiffCSP++[13] as the backbone network of diffusion model, effectively repurposing its exploration to strictly adhere to the symmetry constraints inferred by our method. Specifically, we employ a joint-rectification mechanism on both lattice and fractional coordinates. For the lattice, we constrain its $O(3)$-invariant logarithmic parameterization by applying a binary mask dictated by the inferred crystal family, ensuring the unit cell geometry strictly adheres to the specific constraints on lattice lengths and inter-axial angles defined by the crystal family. Simultaneously, for the fractional coordinates $\mathbf{F}_{t}$, we employ a subspace projection and reconstruction mechanism, effectively rectifying any symmetry violations in each denoising state of fractional coordinates, ensuring that the generated crystal structure strictly resides on the geometric manifold defined by the target Wyckoff symmetries. Further details regarding the lattice masking and coordinate reconstruction are available in Supplementary Information A.4.
@@
Figure 2: Visual comparison of prediction results from DiffCSP++ and our method for the target composition BaP3 ($N=16$). The criteria for stability, uniqueness, and novelty follow the SUN metrics defined by Zeni et al.[30] and are detailed in Supplementary Information C.1.
@@
To evaluate our approach, we benchmark it against a comprehensive suite of generative models. We adopt DiffCSP++ [13] as the primary baseline, as it represents the current state-of-the-art (SOTA) for standard CSP and serves as the most stringent benchmark for composition-conditioned structural generation, particularly for stability, uniqueness, and novelty (SUN) metrics. To provide a broader context for structural reconstruction fidelity, we also encompass CDVAE [25], DiffCSP [12], and CrystaLLM [2] in our comparative analysis, specifically regarding the Matching Rate metric. While MatterGen [30] is also a leading method in materials generation, it is excluded from direct comparison due to a fundamental discrepancy in task formulation. MatterGen employs a joint generation paradigm that co-samples chemical composition and structural coordinates. Although this enables unconstrained exploration, it lacks the strict conditionality inherent to CSP, where the core objective is to resolve the atomic structure for a fixed, pre-specified composition. Thus, a direct quantitative comparison is omitted
@@
We adopt the SUN (Stability, Uniqueness, and Novelty) metrics [30] alongside the Top-20 Matching Rate [12] to provide a comprehensive evaluation of our model. The detailed definitions and calculation protocols for these metrics are provided in Supplementary Information C.1 and C.2. The SUN metrics emphasize physical plausibility, structural diversity, and originality, aspects that are more directly aligned with the objectives of open-ended discovery in materials science; while the Matching Rate prioritizes structural fidelity by quantifying the precision with which the model reconstructs the experimental ground-truth (GT) geometry within defined coordinate and symmetry tolerances. It is worth noting that, these two metrics can be inherently contradictory in a Top-1 evaluation because a model must choose between replicating a known structure to achieve a high Matching Rate or exploring a previously unknown one to achieve high Novelty. This tension is effectively resolved in our Top-20 setting. By generating multiple candidates for each composition, the model is afforded the capacity to simultaneously demonstrate structural fidelity by recovering the experimental ground truth and chemical creativity by proposing stable, novel phases. Thus, reporting both metrics at Top-20 provides a holistic assessment of the model’s dual role as both a precise structural reconstructor and a robust discovery engine for uncharted materials space.
@@
All experiments are conducted on three standard CSP benchmark datasets: Perov-5 [5], MP-20 [9], and MPTS-52 [13]. These datasets were previously used in DiffCSP++ and are employed here with identical training, validation and testing splits to ensure a fair comparison. Detailed descriptions and statistical information regarding these three datasets are available in Supplementary Information C.2. The experimental results for the DiffCSP++ baseline are obtained using its official open-source scripts for standard training and inference [1]. Table 1 presents a side-by-side comparison between the performance of this original method and that of our symmetry-enhanced method.
@@
As shown in Table 1, our method yields notable and consistent improvements over DiffCSP++ across all three benchmarks on the comprehensive SUN metrics. On the MP-20 dataset, our method improves stability and novelty by approximately 124% and 71% respectively relative to the baseline. The advantages are even more pronounced on the challenging MPTS-52 dataset, where our method achieves a relative gain of about 255% in stability and 53% in novelty. Crucially, these component-level improvements compound to yield a striking 376% increase in the overall SUN metric. Even on Perov-5, which features relatively simple structural motifs, our method maintains a clear advantage in stability while matching the saturated uniqueness and novelty scores of the baseline. The significant rise in the overall SUN metric across all benchmark datasets confirms that our method generates candidates that are simultaneously more physically stable and structurally novel, while preserving the high standards of uniqueness.
To tangibly demonstrate the advantage of our symmetry-driven approach over the retrieval-based baseline, we present a comparative case study for the target composition BaP3 ($N=16$) in Fig. 2. As illustrated in Fig. 2a, the DiffCSP++ baseline retrieves a Wyckoff template associated with Space Group 12 ($C2/m$). This configuration assigns Ba atoms to a single $4i$ site and P atoms to a mixed set of $4i$ and $8j$ sites. Such topological mismatch induces severe geometric frustration during structure generation, manifesting as prominent structural distortions with atomic clashes highlighted in red. Consequently, the generated structure is thermodynamically unstable and fails the novelty assessment.
In contrast, Fig. 2b demonstrates that our method successfully infers a symmetry-consistent Wyckoff template ab initio for the predicted space group (No. 14, $P2_{1}/c$). It rigorously assigns atoms to four distinct $4e$ Wyckoff orbits: one for Ba and three for P. By enforcing algebraic consistency, our approach generates a crystal structure fully aligned with the predicted symmetry, yielding a material candidate that satisfies all SUN metrics.
@@
The performance disparity observed in Table  1 and Fig. 2 stems from the fundamental transition from retrieval-based memorization to generative reasoning. The baseline relies on CSPML, which essentially operates as a static lookup system. It rests on the isomorphism assumption, presupposing that for any new material, a structurally identical twin already exists in the training database. Consequently, this approach fails in data-sparse regimes or when exploring truly novel materials where no such prototype exists. When the model forces a novel composition into a retrieved but topologically mismatched template, it introduces geometric frustration. The atoms are coerced into energetically unfavorable coordination environments to satisfy incorrect symmetry constraints, directly leading to poor stability. This reliance on data lookup inherently confines the model to the known structural motifs, rendering it incapable of discovering new polymorphs that lie outside the existing library.
@@
In stark contrast, our method reformulates symmetry determination as ab initio generative inference coupled with rigorous constrained optimization. Instead of memorizing and retrieving existing solutions, our model performs symmetry-aware reasoning to derive valid Wyckoff patterns for unseen compositions strictly from chemical rules. Crucially, a combinatorial optimization algorithm enforces algebraic consistency, ensuring the sum of site multiplicities matches the atomic stoichiometry. This mechanism empowers the model to explore uncharted chemical space while confining stochastic diffusion to a symmetry-compliant configurational space, guaranteeing output rationality. By eliminating symmetry-breaking artifacts of unconstrained generation, our framework enables inverse design of materials with high structural novelty and strict physical rigor.
@@
Complementing the discovery-oriented SUN metrics, Table 2 presents the Top-20 Matching Rate, which quantifies the model’s ability to reconstruct experimentally known ground-truth structures. As shown, our method consistently achieves state-of-the-art performance across all three benchmarks, outperforming the DiffCSP++. While the numerical margin over the previous SOTA is incremental rather than dramatic, this result is scientifically significant when interpreted in the context of generative design. In ab initio materials discovery, the Matching Rate serves primarily as a validation metric rather than the ultimate objective; it confirms that the model’s generative distribution covers the manifold of physically realized structures. A high Matching Rate indicates that the symmetry constraints imposed by our framework do not overly restrict the search space or exclude valid ground-truth configurations. Instead, they effectively guide the diffusion process within a chemically reasonable manifold. Consequently, the combination of superior SUN scores (Table 1) and leading Matching Rates (Table 2) demonstrates that our method resolves the trade-off between exploration and exploitation: it is robust enough to recover known stable phases with high fidelity while possessing the generative flexibility to propose novel, stable polymorphs beyond the training distribution.
@@
In summary, this work develops a symmetry-driven CSP framework that advances beyond conventional retrieval-based baselines. By reducing reliance on pre-existing prototypes, leveraging chemical rules, and enforcing strict algebraic consistency and symmetry compliance, it delivers robust performance in data-sparse regimes and novel material exploration. Experimental results show significant gains over state-of-the-art methods across comprehensive SUN metrics, transcending known structural motifs, as well as superior matching rates. This framework establishes a reliable, highly exploratory pathway for advanced functional material inverse design, redefining CSP from retrieval-based memorization to constrained ab initio generative inference.

### v2

# Overcoming the Combinatorial Bottleneck in Symmetry-Driven Crystal Structure Prediction

@@ Abstract
Crystal structure prediction (CSP), which aims to predict the three-dimensional atomic arrangement of a crystal from its composition, is central to materials discovery and mechanistic understanding. However, given the composition and atomic counts in a unit cell, existing methods struggle with the NP-hard combinatorial challenge of rigorous symmetry enforcement or rely on retrieving known templates, which inherently limits both physical fidelity and the ability to discover genuinely new materials. To solve this, we propose a symmetry-driven generative framework. Our approach leverages large language models to encode chemical semantics and directly generate fine-grained Wyckoff patterns from atomic stoichiometry and counts, effectively circumventing the limitations inherent to database lookups. Crucially, to overcome the exponentially complex problem of combinatorial site assignments, we incorporate domain knowledge through an efficient, linear-complexity heuristic beam search algorithm that rigorously enforces algebraic consistency between site multiplicities and atomic stoichiometry and counts. By integrating this symmetry-consistent template into a diffusion backbone, our approach constrains the stochastic generative trajectory to a physically valid geometric manifold. This framework achieves state-of-the-art performance across stability, uniqueness, and novelty (SUN) benchmarks, alongside superior matching performance, thereby establishing a new paradigm for the rigorous exploration of targeted crystallographic space which can be previously uncharted, with no reliance on existing databases or a priori structural knowledge.
@@
Crystalline materials are fundamental to modern technologies spanning energy, electronics, medicine, and aerospace, and advancing these sectors relies on identifying novel crystal structures with tailored properties [7, 3, 10, 6, 15]. The vastness of the compositional and structural parameter space presents a fundamental bottleneck for materials discovery. Conventional pipelines, guided by chemical intuition and constrained by costly computations and experiments [21, 16], are intrinsically inefficient at exploring this space at scale.
In recent years, artificial intelligence (AI), and deep learning-based generative models in particular, have emerged as a promising route to proposing candidate materials at scale, demonstrating superior generalization capabilities and significant advantages in computational efficiency [11, 2, 13, 29, 31, 8, 30]. Within this landscape, the crystal structure prediction (CSP) task [11, 13, 8], which seeks the three-dimensional arrangement of atoms in the unit cell from a given composition, plays a pivotal role in AI-driven materials design workflows [29]. Diffusion models have become the dominant architecture for 3D structure generation [27, 24], and, for the CSP task, diffusion-based approaches also deliver state-of-the-art performance [13, 30].
However, the CSP task is fundamentally governed by stringent crystallographic rules, imposing rigid constraints on the design methodology. Physically plausible crystal structures typically obey the symmetry constraints imposed by specific space groups, which strictly constrain the allowed Wyckoff positions, determining atomic multiplicities, site symmetries, and the relative arrangement of atoms within the unit cell. Only when the given chemical composition is mathematically consistent with the corresponding Wyckoff positions can physically reasonable crystal structures be obtained.
@@
However, rigorously enforcing these physical constraints imposed by space groups remains a universal bottleneck across the entire CSP domain. Fundamentally, determining an exact combination of Wyckoff positions whose multiplicities perfectly align with the exact atom counts of a given composition in the unit cell constitutes an inherently NP-hard combinatorial optimization problem, characterized by an exponentially exploding search space. Current methods struggle to efficiently resolve this task, often resorting to partial compromises. For example, aiming at efficiently addressing site assignments, Xin et al. [26] introduces a search method for Wyckoff positions with a pruning strategy; however, the approach remains a brute-force enumeration that inevitably faces this intractable exponential complexity for large atomic systems, and it lacks a likelihood-based ranking for the generated candidates.
Within the landscape of generative AI, existing methods also only partially circumvent this challenge. Early and generic 3D generative models, such as CDVAE [25] and DiffCSP [12], completely bypass this discrete combinatorial challenge. They operate directly on continuous coordinates and rely on neural networks to implicitly learn crystallographic symmetries, which leaves the generative trajectory unconstrained in an unimaginably vast search space. CrystaLLM [2] presents a text generation framework autoregressively predicting Crystallographic Information File (CIF) formats; however, any generated symmetry or Wyckoff symbols are mere textual probabilistic outputs, entirely lacking algebraic or physical hard constraints. Some recent approaches condition the generator on a space group label [29, 30] but do not explicitly represent or enforce Wyckoff positions, so symmetry is only captured in a coarse global sense.
Moving beyond coarse global symmetry, several recent AI methods attempt to explicitly utilize and strictly enforce Wyckoff positions, yet they remain hindered by the intractable complexity of exact site assignments. While WyCryst [31] elegantly guides generation via Wyckoff-based latent representations, its soft-penalty sampling cannot guarantee strict stoichiometric alignment. Consequently, extracting 3D structures consistent with Wyckoff positions relies on post-hoc filtering, a process that can become computationally prohibitive given the NP-hard nature of combinatorial site assignments. CrystalFormer [4] presents a compelling framework representing crystals as Wyckoff sequences with logit masking for symmetry enforcement. While effective for generating structures within a given space group number, the final stoichiometry is an emergent outcome of the sampling process rather than a hard constraint. Since the model focuses on generative modeling, it does not employ an explicit constrained optimization mechanism to align site multiplicities with a given composition and atomic number in a cell. DiffCSP++ [13] goes further by enforcing strict space group symmetry defined by Wyckoff templates during the 3D material generation process. It employs two strategies for obtaining these templates: the first directly utilizes the ground-truth Wyckoff templates of the target material to guide structure generation. However, in practical CSP scenarios, these ground-truth templates are inherently unknown a priori. As an alternative, it relies on a retrieval-based approach, which assumes that a suitable space group and Wyckoff template are already known in the given lookup database. In practice, these templates are retrieved from existing structures via metric learning methods, such as CSPML [14]. However, restricting generation to retrieved templates confines the model to known symmetry patterns, thereby limiting its ability to discover genuinely new space-group and Wyckoff configurations for a given composition.
@@
To address these fundamental limitations, we propose an efficient symmetry-driven generative framework, which enables the ab initio generation of fine-grained Wyckoff site assignments directly from composition and atom counts, independent of any structural priors, as depicted in Fig. 1. Building upon the probabilistic distributions inferred by large language models (LLMs), we explicitly overcome the intractable exponential complexity of combinatorial site assignments discussed above by designing an efficient heuristic constrained-optimization search. This algorithm drastically compresses the time complexity from an exponential scale down to a linear scale, rendering the search computationally tractable while rigorously enforcing algebraic consistency between site multiplicities and atomic stoichiometry without relying on computationally prohibitive brute-force enumeration. Crucially, the predicted symmetries are then imposed as hard geometric constraints to guide and modulate the diffusion process for three-dimensional crystal geometries, thereby rectifying the denoising trajectory, confining the generative process to a physically valid manifold, guaranteeing both structural plausibility and computational precision. As a result, this leads to a significant enhancement in the stability, uniqueness, and novelty (SUN) [30] of the predicted materials, as well as the matching rate metrics [12].
@@
Below, we outline the core methodology, with full implementation details, as well as training and testing configurations provided in Appendix A and B, respectively.
@@
The second LLM, denoted as $LLM_{w}$, takes the same expanded atomic sequence together with the predicted space group $S_{g}$ and outputs a probability distribution over Wyckoff letters compatible with that space group for each atom. Rather than concatenating $S_{g}$ naively to the input, we encode the predicted space group into a continuous embedding and inject it into each transformer block via the feature-wise linear modulation (FiLM) [18] module. This FiLM-based conditioning ensures that Wyckoff letter predictions are explicitly modulated by the inferred space group, rather than being learned in an implicitly entangled fashion. Concretely, FiLM($S_{g}$) produces per-channel scaling and shifting coefficients that modulate the intermediate representations of $LLM_{w}$, allowing the model to adapt its representation to the symmetry class:
$p_{w}(l_{n}=l|C_{\text{atoms}})=LLM_{w}\big(C_{\text{atoms}},\mathrm{FiLM}(S_{g})\big),$ (2)
@@
where $p_{w}(l_{n}=l|C_{\text{atoms}})$ denotes the probability that the $n$-th atom is assigned to letter $l$, $L(S_{g})$ is the set of admissible Wyckoff letters for space group $S_{g}$. Specific details regarding the network architectures and hyperparameter configurations of these two large language models are provided in Appendix A.1-A.2 and B.
Second, we formulate the Wyckoff letter assignment for each atom as a constrained optimization problem based on the probability distribution $p_{w}(l_{n}|C_{\text{atoms}})$ predicted by the LLM, and propose an efficient algorithm to solve it. This formulation strictly adheres to the combinatorial constraints imposed by the Wyckoff site multiplicity rules under the inferred space group symmetry. Specifically, the constrained maximization objective is defined as follows:
@@
where $\text{mult}(l_{n})$ refers to the multiplicity of Wyckoff letter $l_{n}$, and $\text{Count}_{e}(l_{n})$ is the number of atoms of element $e$ assigned to $l_{n}$, $N_{e}$ is the number of atoms of element $e$ in a cell, the set $\{l_{n}^{*}\}_{n=1}^{N}$ represents the optimal Wyckoff letter assignment. The key point here is that $\text{Count}_{e}(l_{n})$ must be an integer multiple of $\text{mult}(l_{n})$, a hard constraint that cannot be satisfied by simply using the maximum probability outputs of each token from the LLM. Fundamentally, deriving the exact global optimum under these precise constraints constitutes an NP-hard combinatorial optimization problem. Consequently, finding the solution for Eq. (3) via brute-force enumeration is computationally intractable, as the search space grows exponentially with the number of atoms, scaling as $O(|L(S_{g})|^{N})$. To overcome this bottleneck, we employ propose a beam search algorithm to efficiently solve the optimization problem. By maintaining only the most promising candidates within a fixed beam width, this strategy drastically compresses the time complexity from exponential to a linear $O(N)$, rendering the prediction of physically valid Wyckoff assignments computationally tractable. Comprehensive details regarding this algorithm are provided in Appendix A.3.
Third, we guide and constrain the diffusion-based generation of 3D atomic structures using the predicted space group and Wyckoff site template. To instantiate this approach, we adopt the graph neural network (GNN) applied by DiffCSP++[13] as the backbone network of diffusion model, effectively repurposing its exploration to strictly adhere to the symmetry constraints inferred by our method. Specifically, we employ a joint-rectification mechanism on both lattice and fractional coordinates. For the lattice, we constrain its $O(3)$-invariant logarithmic parameterization by applying a binary mask dictated by the inferred crystal family, ensuring the unit cell geometry strictly adheres to the specific constraints on lattice lengths and inter-axial angles defined by the crystal family. Simultaneously, for the fractional coordinates $\mathbf{F}_{t}$, we employ a subspace projection and reconstruction mechanism, effectively rectifying any symmetry violations in each denoising state of fractional coordinates, ensuring that the generated crystal structure strictly resides on the geometric manifold defined by the target Wyckoff templates. Further details regarding the lattice masking and coordinate reconstruction are available in Appendix A.4.
@@
Figure 2: Visual comparison of prediction results from DiffCSP++ and our method for a testing sample, i.e., BaP3 ($N=16$). The definition for Stability, Uniqueness, and Novelty are detailed in Appendix C.1.
(a) Retrieved and Prediction results of DiffCSP++.
(b) Prediction results of our method.
Figure 3: Visual comparison of prediction results from DiffCSP++ and our method for a testing sample, i.e., K2AgBiI6 ($N=40$). The definition for Stability, Uniqueness, and Novelty are detailed in Appendix C.1.
(a) Retrieved and Prediction results of DiffCSP++.
(b) Prediction results of our method.
Figure 4: Visual comparison of prediction results from DiffCSP++ and our method for a testing sample, i.e., Lu2AlRu ($N=16$). The definition for Stability, Uniqueness, and Novelty are detailed in Appendix C.1.
@@
To evaluate our approach, we benchmark it against a comprehensive suite of generative models. We adopt DiffCSP++ [13] as the primary baseline, as it represents the current state-of-the-art (SOTA) for standard CSP and serves as the most stringent benchmark for composition-conditioned structural generation. To provide a broader context for structural reconstruction fidelity, we also encompass CDVAE [25], DiffCSP [12], and CrystaLLM [2] in our comparative analysis, specifically regarding the Matching Rate metric. While MatterGen [30] and CrystalFormer [4] are also leading methods in materials generation, they are excluded from direct comparison due to a fundamental discrepancy in task formulation. They employ a joint generation paradigm that co-samples chemical composition and structural coordinates. While this enables broad exploration of the joint space of compositions and structures, it lacks the strict conditionality inherent to CSP, where the core objective is to resolve the atomic structure for a fixed, pre-specified composition. Thus, a direct quantitative comparison is omitted.
@@
To provide a comprehensive evaluation of our model, we adopt the Top-20 SUN (Stability, Uniqueness, and Novelty) metrics based on the core principles of [30], alongside the Top-20 Matching Rate [12]. The detailed definitions and adapted calculation protocols for these metrics, designed to accommodate our extensive Top-20 evaluation setting, are provided in Appendices C.1 and C.2. The SUN metrics emphasize physical plausibility, structural diversity, and originality, aspects that are more directly aligned with the objectives of open-ended discovery in materials science; while the Matching Rate prioritizes structural fidelity by quantifying the precision with which the model reconstructs the experimental ground-truth (GT) geometry within defined coordinate and symmetry tolerances. It is worth noting that, these two metrics can be inherently contradictory in a Top-1 evaluation because a model must choose between replicating a known structure to achieve a high Matching Rate or exploring a previously unknown one to achieve high Novelty. This tension is effectively resolved in our Top-20 setting. By generating multiple candidates for each composition, the model is afforded the capacity to simultaneously demonstrate structural fidelity by recovering the experimental ground truth and chemical creativity by proposing stable, novel phases. Thus, reporting both metrics at Top-20 provides a holistic assessment of the model’s dual role as both a precise structural reconstructor and a robust discovery engine for uncharted materials space.
@@
All experiments are conducted on three standard CSP benchmark datasets: Perov-5 [5], MP-20 [9], and MPTS-52 [13]. These datasets were previously used in DiffCSP++ and are employed here with identical training, validation and testing splits to ensure a fair comparison. Detailed descriptions and statistical information regarding these three datasets are available in Appendix C.2. The experimental results for the DiffCSP++ baseline are obtained using its official open-source scripts for standard training and inference [1]. Table 1 presents a side-by-side comparison between the performance of this original method and that of our symmetry-enhanced method.
@@
As shown in Table 1, our method yields notable and consistent improvements over DiffCSP++ across all three benchmarks on the comprehensive SUN metrics. On the MP-20 dataset, our method improves stability and novelty by approximately 124% and 71% respectively relative to the baseline. The advantages are even more pronounced on the challenging MPTS-52 dataset, where our method achieves a relative gain of about 255% in stability and 53% in novelty. Crucially, these component-level improvements compound to yield a striking 376% increase in the overall SUN metric. Even on Perov-5, which features relatively simple structural motifs, our method maintains a clear advantage in stability while matching the saturated uniqueness and novelty scores of the baseline. The significant rise in the overall SUN metric across all benchmark datasets confirms that our method generates candidates that are simultaneously more physically stable and structurally novel, while preserving the high standards of uniqueness. Figs. 2-4 visualize the prediction results for representative testing samples: BaP3, K2AgBiI6, and Lu2AlRu. The baseline DiffCSP++ forces these compositions into retrieved, topologically mismatched Wyckoff templates, inducing prominent atomic clashes. Consequently, these baseline structures fail both the stability and novelty criteria. In contrast, our approach infers compatible space groups and Wyckoff assignments directly from the stoichiometry. This resolves the structural conflicts and generates candidate materials that successfully satisfy all three criteria: Stability, Uniqueness, and Novelty.
@@
The performance disparity observed in Table  1 and Fig. 2-4 stems from the fundamental transition from retrieval-based memorization to generative reasoning. The baseline relies on CSPML, which essentially operates as a static lookup system. It rests on the isomorphism assumption, presupposing that for any new material, a structurally identical twin already exists in the training database. Consequently, this approach fails in data-sparse regimes or when exploring truly novel materials where no such prototype exists. When the model forces a new querying composition into a retrieved but topologically mismatched template, it introduces geometric frustration. The atoms are coerced into energetically unfavorable coordination environments to satisfy incorrect symmetry constraints, directly leading to poor stability. This reliance on data lookup inherently confines the model to the known structural motifs, rendering it incapable of discovering new polymorphs that lie outside the existing library.
@@
In stark contrast, our method reformulates symmetry determination as ab initio generative inference coupled with rigorous constrained optimizationInstead of retrieving existing solutions, the model learns a unified representation space from diverse structural data, enabling it to generalize and derive valid Wyckoff patterns for unseen compositions based on chemical rules. Crucially, a heuristic combinatorial optimization algorithm enforces algebraic consistency, ensuring the sum of site multiplicities matches the atomic stoichiometry. This mechanism empowers the model to explore uncharted chemical space while confining stochastic diffusion to a symmetry-compliant configurational space, guaranteeing output rationality. By eliminating symmetry-breaking artifacts of unconstrained generation, our framework enables inverse design of materials with high structural novelty and strict physical rigor.
@@
Complementing the discovery-oriented SUN metrics, Table 2 presents the Top-20 Matching Rate, which quantifies the model’s ability to reconstruct experimentally known ground-truth structures. As shown, our method consistently achieves state-of-the-art performance across all three benchmarks, outperforming all of the baseline methods. This result is scientifically significant when interpreted in the context of generative design. In ab initio materials discovery, the Matching Rate confirms that the model’s generative distribution covers the manifold of physically realized structures. A high Matching Rate indicates that the symmetry constraints imposed by our framework do not overly restrict the search space or exclude valid ground-truth configurations. Instead, they effectively guide the diffusion process within a chemically reasonable manifold. Consequently, the combination of superior SUN scores (Table 1) and leading Matching Rates (Table 2) demonstrates that our method resolves the trade-off between exploration and exploitation: it is robust enough to recover known stable phases with high fidelity while possessing the generative flexibility to propose novel, stable polymorphs beyond the existing samples.
@@
In summary, this work presents a symmetry-driven CSP framework that advances beyond conventional paradigm. By reducing reliance on pre-existing prototypes, leveraging chemical rules, and enforcing strict algebraic consistency and symmetry compliance through an efficient generative framework incorporating probability distribution predictions from large language models and an efficient heuristic search, it renders an inherently NP-hard combinatorial optimization problem computationally tractable. Consequently, it delivers robust performance in data-sparse regimes and novel material exploration. Experimental results show significant gains over state-of-the-art methods across comprehensive SUN metrics, transcending known structural motifs, as well as superior matching rates. This framework establishes a reliable, highly exploratory pathway for advanced functional material inverse design, redefining CSP from retrieval-based memorization to constrained ab initio generative inference.